OpenStudy (anonymous):

(2+square root -3/2)(2-square root -3/2)

OpenStudy (anonymous):

Again... \[\sqrt{-1}=i\]

Parth (parthkohli):

Difference of two squares ---> \((a + b)( a- b) = a^2 - b^2\) Can you do it from now?

OpenStudy (anonymous):

\[\left(2+\frac{\sqrt{-3}}{2}\right)\left(2-\frac{\sqrt{-3}}{2}\right)\] like that?

OpenStudy (anonymous):

or \[(\frac{2+\sqrt{-3}}{2})(\frac{2-\sqrt{-3}}{2})\]